OpenStudy (anonymous):

X walks down an UP-escalator and counts 150 steps.Y walks up the same same escalator and count 75 steps.X takes three times as many steps in a given time as Y.How many steps are visible on the escalator?

OpenStudy (anonymous):

When time is constant if X = 150 steps ,Y = 150/3 = 50 steps and Escalator assume takes E steps. For Y to take 75 steps Escalator will move 1.5E 150 -E = 75 +1.5 E => E = 30 Ans 150 -30 = 120

OpenStudy (anonymous):

am i correct???

OpenStudy (lalaly):

let T be time Y takes to make 25 steps. so Y takes 3T to make 75, and X takes 2T to make 150. Suppose the escalator has N steps and moves A steps in time T so X covers N + 2A = 150 N - 3A = 75. Hence N = 120, A= 15.

OpenStudy (anonymous):

i get the answer as 120 steps. tell me if anybody else gets the same answer.

OpenStudy (anonymous):

Your mind works so beautifully...Great yaar

OpenStudy (anonymous):

that seems to be what everyone is getting

OpenStudy (lalaly):

lol thanx:D